A STUDY OF THE INTERMOLECULAR ν_{5}^{1} VIBRATION IN OC-H^{35} Cl BASED ON NEAR INFRARED SPECTROSCOPY

R. GARNICA, A. I. MCINTOSH, Z. WANG, R. R. LUCCHESE, \underline{J. W. BEVAN}, Department of Chemistry, Texas A&M University, College Station, Texas 77843-3255; A. R. W. MCKELLAR, Steacie Institute, National Research Council, Ottawa, Ontario K1A OR6, Canada.